You take the bus to school 5 days per week. The waiting time for the next bus to arrive at your bus stop is normally distributed with a mean of 20 minutes and standard deviation of 5 minutes. Assume that the waiting times are independent from day to day. What is the probability that you wait more than 15 minutes for your bus to arrive on 3 of the 5 days next week?
Answer by stanbon(75887)

You take the bus to school 5 days per week. The waiting time for the next bus to arrive at your bus stop is normally distributed with a mean of 20 minutes and standard deviation of 5 minutes. Assume that the waiting times are independent from day to day. What is the probability that you wait more than 15 minutes for your bus to arrive on 3 of the 5 days next week?
--------
Find probability of >= 15 minute wait:
z(15) = (15-20)/5 = -1
P(x > 15) = P(z > -1) = 0.8413
-----
Find probability of 3 in 5.
Binomial Problem with n = 5 and p = 0.8413
--
P(x = 3) = 5C3(0.8413)^3(0.1587)^2 = 0.1500
=================================================
Cheers,
Stan H.